OpenStudy (anonymous):

Medal to whoever helps. (Algebra 1)

OpenStudy (anonymous):

what do you need help with

OpenStudy (adamaero):

help

geerky42 (geerky42):

Algebra 1, duh.

OpenStudy (anonymous):

The following function defines a recursive sequence: f(0) = -5 f(1) = 20 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -5, -20, -65, -200, … -5, 20, -92, 372, … -5, -24, -92, -372, … -5, 20, -65, 200, …

OpenStudy (adamaero):

answer: ☺

OpenStudy (anonymous):

I can't see anything...

geerky42 (geerky42):

Try refresh page.

OpenStudy (anonymous):

can you explain it to me? The problem...

OpenStudy (anonymous):

please help.

geerky42 (geerky42):

So you are given that \(f(0) = -5\) and \(f(1)=20\). we need to find next few terms: \(f(2),~f(3),~\ldots\) we are given equation \(f(n) = -4~f(n -1) - 3~f(n - 2)\) So to find \(f(2)\), you have \(f(\color{red}{2}) = -4~f(\color{red}{2} -1) - 3~f(\color{red}{2} - 2)\\~~~~~~~=-4~f(1) - 3~f(0)\) We are given that \(f(0) = -5\) and \(f(1)=20\), right? So \(\cdots=-4 (20)-3(-5)=\cdots~?\)

geerky42 (geerky42):

Is that clear?

OpenStudy (anonymous):

0.o

geerky42 (geerky42):

Doesn't sound good...

OpenStudy (anonymous):

ok so \[n\] equals what ever number we need?

geerky42 (geerky42):

Right.

OpenStudy (anonymous):

I'm starting to get it.

geerky42 (geerky42):

ok, so can you find third term: \(f(2)\)?

OpenStudy (anonymous):

hang on...

OpenStudy (anonymous):

\[f(2)=-4f(1)-34\] ?

OpenStudy (anonymous):

that last bit was supposed to be 3f

geerky42 (geerky42):

\(f(2)=-4f(1)-3f(0)\)

OpenStudy (anonymous):

alright. now what?

geerky42 (geerky42):

You just replace \(f(1)\) with \(20\) and \(f(0)\) with \(-5\) So you have \(f(2)=-4[20]-3[-5]\)

OpenStudy (anonymous):

so it should be b?

geerky42 (geerky42):

A or D to me? You got -92?

OpenStudy (anonymous):

-95

geerky42 (geerky42):

\(-4[20]=-80\) and \(-3[-5]=15\) So \(f(2)=-80+15=\cdots~?\)

OpenStudy (anonymous):

oh you add them?

OpenStudy (anonymous):

wait that still doesn't work...

geerky42 (geerky42):

well, that's because you multiply negative number. Made it positive, so you add them.

OpenStudy (anonymous):

it adds up to -65

geerky42 (geerky42):

Right, so that leave us A or D. Now find \(f(3)\). From recursive formula, we have \(f(3) = -4f(2)-3f(1)\) Replace \(f(2)\) with \(-65\) and \(f(1)\) with \(20\)

OpenStudy (anonymous):

but earlier you said -92...

geerky42 (geerky42):

No? I just asked if you got -92? f(2) is actually -65.

OpenStudy (anonymous):

oh, so f(3) is 320?

geerky42 (geerky42):

No? You have \(f(3)=-4[-65]-3[20]\), right? -4 times -65 is 260, right? -3 times 20 is -60, right? So you have \(f(3)=260-60=\cdots~?\)

OpenStudy (anonymous):

I ended up with \[f(3)= 260+60\] because \[-3(20)\] is a negative right?

geerky42 (geerky42):

Yeah it is negative, so you are supposed to subtract 60 from 260

geerky42 (geerky42):

if you have two minus signs, they will cancel each other out, so you end up with positive. But this time, you only have one minus sign. so that's subtraction.

OpenStudy (anonymous):

but isn't that a double negative (sorry about my stupidity I'm awful at math)

geerky42 (geerky42):

You multiplied -3 by 20, which is -60. 20 is positive.

OpenStudy (anonymous):

ok.

geerky42 (geerky42):

So does it make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so D.?

geerky42 (geerky42):

Ok, so your answer is D. I didn't notice that 20 is negative in A, lol...

OpenStudy (anonymous):

Thank you SOOOOO much!

geerky42 (geerky42):

No problem